OpenStudy (anonymous):

A plane is flying at an elevation of 20000 feet. It is within sight of the airport and the pilot finds that the angle of depression to the airport is 23 degree. What is the distance between the plane and the airport?

OpenStudy (anonymous):

Cos(degree) = adjacent/hypotenuse cos(23) = 20,000/x x = 20,000/cos(23) x = 21727.2 feet